OpenStudy (anonymous):

PLEASE I BEG YOU, HELP ME, I'M DESPARTE What is the solution to the equation 2(3^x+1)+8=108???

OpenStudy (anonymous):

i'll try :) is it like this \(\sf 2(3^x+1)+8=108 \) or \(\sf 2(3^{x+1})+8=108 \)

OpenStudy (anonymous):

yes that's the quesion

OpenStudy (misty1212):

\[2\times 3^{x+1}+8=108\] subtract \(8\) first and you have \[2\times 3^{x+1}=100\] then divide by 2 and get \[3^{x+1}=50\]

OpenStudy (misty1212):

from there it depends on what kind of answer you are looking for if you want a numerical answer, use the change of base formula \[x+1=\frac{\ln(50)}{\ln(3)}\] then subtract 1 and use a calculator

OpenStudy (anonymous):

omg, thank-you so very much

OpenStudy (anonymous):

wait hold on, actually why did you put ln 50 over ln 3???

OpenStudy (misty1212):

yw the final answer will be \[x=\frac{\ln(50)}{\ln(3)}-1\]

OpenStudy (anonymous):

why not just 50/3???

OpenStudy (misty1212):

ok lets go slow a little

OpenStudy (misty1212):

this is called the "change of base" formula \[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (misty1212):

you have \[3^{x+1}=50\] so \[x+1=\frac{\ln(50)}{\ln(3)}\]

OpenStudy (misty1212):

clear that up a little? i can show you another example if you like

OpenStudy (anonymous):

no it's fine, thanks

OpenStudy (misty1212):

lets go ahead and find the number http://www.wolframalpha.com/input/?i=ln%2850%29%2Fln%283%29-1

OpenStudy (anonymous):

actually that's all there is, but thanks

OpenStudy (misty1212):

of course since we are going to end up with a calculator anyways, lets see if the answer is right http://www.wolframalpha.com/input/?i=2*3^%28x%2B1%29%2B8%3D108

OpenStudy (anonymous):

i'm just confused.. the question says: 2(3^x+1)+8=108, not 2*3^(x+1)+8=108 ? so it means it should be \(\sf 2(3^x+1)+8=108\\2(3^x+1)=100\\3^x+1=50\\3^x=49\) then apply the change of base...? @misty1212